# Matching-Based Policy Learning

### v1

@@ Abstract
Treatment heterogeneity is ubiquitous in many areas, motivating practitioners to search for the optimal policy that maximizes the expected outcome based on individualized characteristics. However, most existing policy learning methods rely on weighting-based approaches, which may suffer from high instability in observational studies. To enhance the robustness of the estimated policy, we propose a matching-based estimator of the policy improvement upon a randomized baseline. After correcting the conditional bias, we learn the optimal policy by maximizing the estimate over a policy class. We derive a non-asymptotic high probability bound for the regret of the learned policy and show that the convergence rate is almost $1/\sqrt{n}$. The competitive finite sample performance of the proposed method is demonstrated in extensive simulation studies and a real data application.
Keywords: individualized treatment rules, matching, observational study, policy learning, regret bound
@@ 1 Introduction
Recently, data-driven personalized decision-making has received increased attention across various fields. In healthcare, clinicians utilize patients’ characteristics and medical history to tailor diabetes management, aiming to enhance health outcomes (Bertsimas et al., 2017). In public policy, leveraging individualized features can lead to an improved allocation of social services and thus enhance overall efficiency (Kube et al., 2019). In product recommendations, the online platforms can effectively keep a significant number of customers engaged by recommending the products that match their preferences (Bastani et al., 2022). A common feature underlying these applications is the presence of treatment heterogeneity, which is a blessing that brings opportunities to improve the expected outcome in the target population. Therefore, it is imperative to exploit such heterogeneity efficiently and learn the optimal policy that allocates treatments based on individualized characteristics.
@@
Statistical methods for estimating the optimal policy have been extensively developed in the literature, among which can be roughly divided into model-based methods and direct-search methods (e.g., Qian and Murphy, 2011; Zhao et al., 2012; Zhou et al., 2017; Kitagawa and Tetenov, 2018; Shi et al., 2018). For instance, Q-learning (Qian and Murphy, 2011) and A-learning (Shi et al., 2018) first fit the mean outcome given covariates and treatment, or posit a model for the contrast function, then determine the optimal policy based on the corresponding estimate. Both methods belong to model-based approaches, and their effectiveness relies heavily on correct model specification. Alternatively, outcome weighted learning (O-learning, Zhao et al., 2012) is one of the pioneer works of direct-search methods that nonparametrically learns the optimal policy by utilizing the inverse probability weighting estimator (IPWE) for the mean outcome corresponding to the policy, i.e., value function. In observational studies, the unknown propensity score in O-learning needs to be estimated using logistic regression or machine learning approaches (Lee et al., 2010) for adjusting confounding bias. One recognized potential drawback of weighting-based approaches is the instability caused by extreme estimated propensity scores or model misspecification.
@@
To provide additional robustness, some policy learning literature based on augmented IPWE (AIPWE) of value function has emerged (e.g., Zhang et al., 2012a, b; Zhao et al., 2019; Athey and Wager, 2021; Pan and Zhao, 2021), which combines the strengths of both the aforementioned strategies. Zhang et al. (2012b) proposed an AIPWE of value function that possesses doubly robustness in the sense that the estimator is consistent if either propensity score model or outcome mean model is correctly posited, where parametric models were used for these nuisance functions. Yet, the performance could be unstable when both models are misspecified (Kang and Schafer, 2007). The seminal work by Athey and Wager (2021) leverages cross-fitting technique (Chernozhukov et al., 2018) and utilizes flexible machine learning approaches for nuisance functions. They demonstrated that the learned policy achieves minimax optimal regret if the mean-squared errors of nuisance functions satisfy some order conditions of convergence rate. Though the promising theoretical guarantee, this AIPWE-based method still suffers from extreme weights, especially when the sample size is relatively small. Similar phenomena have been discussed in both policy learning and causal inference literature (e.g., Zhao et al., 2019; Wu et al., 2022). Hence, it is crucial to develop a robust approach for policy learning in observational studies with favorable finite sample performance.
@@
In this paper, we propose a matching-based method (MB-learning) for learning the optimal policy with observational data. Matching is an appealing approach for estimating causal effect in observational studies, which balances covariates between different groups intuitively and exhibits robustness against extreme weights. There is a significant amount of work on matching estimators (see, e.g., Imbens and Rubin (2015) and Rosenbaum (2020) for literature review). The insightful idea behind MB-learning originates from an alternative form of the value function based on the pair of potential outcomes (referred to as advantage function in Section 2), which can be interpreted as the value function improvement upon a randomized baseline. Motivated by the advantage function, we impute the missing counterfactual outcomes to develop the corresponding estimator, where we utilize nearest-neighbor (NN) matching on covariates (Abadie and Imbens, 2006). One recognized limitation of matching estimator is its susceptibility to the curse of dimensionality, as it is not $\sqrt{n}$-consistent in general due to the conditional bias. To mitigate this issue, we employ a bias correction technique proposed by Abadie and Imbens (2011) and develop an improved matching-based estimator with $\sqrt{n}$-asymptotic normality. Leveraging the enhanced estimator, we recast the policy learning optimization into a weighted classification problem and derive the optimal policy.
This paper makes four major contributions to existing literature:
  1. 1.
    Distinct from the (A)IPWE-based approaches, MB-learning is a robust method for learning the optimal policy in observational studies. By utilizing a bias-corrected matching-based estimator, MB-learning mitigates the potential variability arising from propensity score model misspecification or extreme weights.
  2. 2.
    We establish theoretical guarantee for the optimal policy learned by the matching-based estimator, where a non-asymptotic high probability bound for the value function difference between the best-in-class policy and the learned policy is derived.
  3. 3.
    We demonstrate the superiority of MB-learning both theoretically and empirically. Specifically, Athey and Wager (2021) established that their approach achieves rate-optimality in terms of minimax regret, which is on the order of $1/\sqrt{n}$. In contrast, the regret bound of MB-learning is almost $1/\sqrt{n}$. However, as illustrated in simulation studies and real data analysis, MB-learning exhibits more robust finite sample performance.
  4. 4.
    To our knowledge, matching has rarely been utilized for policy learning. The only exception is matched learning (Wu et al., 2020, M-learning), where they estimated the optimal policy by weighted support vector machine and presented the competitive performance in observational studies. We highlight several advantages of MB-learning. First, MB-learning is more intuitive that directly targets an estimator of the advantage function, while M-learning seeks to maximize an empirical analogue of a complicated objective function (Wu et al., 2020, Theorem 3.1). Second, the policy derived from MB-learning possesses appealing interpretability and implementability for practitioners, while that from M-learning with Gaussian kernel lacks of interpretation. Third, the presence of bias introduced by NN matching could substantially deteriorate the performance of M-learning when dealing with moderately high-dimensional covariates.
@@
The rest of this paper is organized as follows. In Section 2, we develop the methodology of MB-learning. Specifically, we formulate the treatment assignment framework and introduce the AIPWE-based policy learning approach developed by Athey and Wager (2021). Then we propose a matching-based estimator of the advantage function, which is further improved by a bias correction technique. Finally, the optimal policy is derived from the resulting estimate. In Section 3, we establish rigorous theoretical guarantee for MB-learning. The empirical performance of MB-learning is demonstrated through extensive simulation studies in Section 4, while the application to National Supported Work (NSW) Program is presented in Section 5. We conclude this paper with discussion in Section 6.
@@ 2.1 Notations and Preliminaries
[…]
@@
We consider the problem of learning a policy $\pi$, a decision function from covariates space $\mathcal{X}$ to treatment space $\{0,1\}$, with observational data. In practice, it is unrealistic to assign a complex policy to the target population. We should attach problem-specific constraints to our learned policy, such as political restrictions (Kitagawa and Tetenov, 2018), fairness (Athey and Wager, 2021) and interpretability (Zhang et al., 2012a). Therefore, we seek to learn the optimal policy within a restricted policy class $\Pi$. The value function $V(\pi)$ relative to $\pi$ (Zhao et al., 2012) is the expected outcome after implementing the treatment allocation by $\pi$, defined as $V(\pi)=E\left[Y(\pi(X))\right]$. The optimal policy $\pi^{*}$ within class $\Pi$ maximizes the value function $V(\pi)$, i.e., $\pi^{*}=\arg\max_{\pi\in\Pi}V(\pi)$.
@@
The advantage function serves as an alternative form of the value function:
$A(\pi)=E\left[\left(2\pi(X)-1\right)\tau(X)\right].$
@@
Since $V(\pi)=E[Y(0)+(Y(1)-Y(0))\pi(X)]$, it follows that $A(\pi)=2\left[V(\pi)-\frac{E[Y(1)]+E[Y(0)]}{2}\right]$ by algebraic manipulations. Hence, the advantage function $A(\pi)$ can be interpreted as the improvement achieved by implementing policy $\pi$ relative to a randomly assigned policy with equal probability (Mo et al., 2021). Moreover, the policy that maximizes $A(\pi)$ is equivalent to that maximizes $V(\pi)$. Throughout this paper, we focus on policy learning based on $A(\pi)$.
@@
Following the idea of semiparametric efficient estimation (Robins et al., 1994), some existing literature developed AIPWEs of $A(\pi)$ and subsequently determined the optimal policies based on these estimators (e.g., Zhang et al., 2012a; Athey and Wager, 2021). Notably, the AIPWE proposed by Athey and Wager (2021) is $\hat{A}_{aipw}(\pi)=n^{-1}\sum_{i=1}^{n}\left(2\pi(X_{i})-1\right)\Gamma_{i}$, where
$\Gamma_{i}=\mu(X_{i},1)-\mu(X_{i},0)+\frac{W_{i}-e(X_{i})}{e(X_{i})\left(1-e(X_{i})\right)}\left(Y_{i}-\mu(X_{i},W_{i})\right).$
The advantage of this AIPWE for $A(\pi)$ is its doubly robustness, providing additional protection from model misspecification in observational studies. Athey and Wager (2021) employed black-box machine learning methods, which achieve sufficiently fast convergence rate in mean-squared error, to estimate the nuisance functions. They developed a plug-in estimator of $\hat{A}_{aipw}(\pi)$ and learned the optimal policy based on the resulting estimator. However, as previously mentioned, estimating the unknown propensity score $e(x)$ in the denominator of $\Gamma_{i}$ may introduce instability, particularly when the estimated propensity score is extreme. In addition, the estimation for nuisance functions by machine learning methods may be unsatisfactory with a relatively small sample size, even though the performance can be guaranteed in large sample. Therefore, it is imperative to develop a more sensible approach for learning the optimal policy in observational studies, while ensuring desirable efficiency and robustness in finite sample.

### v2

@@ Abstract
The beneficial effects of treatments vary across individuals in most studies. Treatment heterogeneity motivates practitioners to search for the optimal policy based on personal characteristics. A long-standing common practice in policy learning has been estimating and maximizing the value function using weighting techniques. Matching is widely used in many applied disciplines to infer causal effects, which is intuitively appealing because the observed covariates are directly balanced across different treatment groups. Nevertheless, matching is rarely explored in policy learning. In this work, we propose a matching-based policy learning framework. We adapt standard and bias-corrected matching methods to estimate an alternative form of the value function: the advantage function, which can be interpreted as the expected improvement achieved by implementing a given policy compared to the equiprobable random policy. We then learn the optimal policy over a restricted policy class by maximizing the matching estimator of the advantage function. We derive a non-asymptotic high probability bound for the regret of the learned optimal policy. Moreover, we show that the learned policy is almost rate-optimal. The competitive finite sample performance of the proposed method compared to weighting-based and outcome modeling-based learning methods is demonstrated in extensive simulation studies and a real data application.
Keywords: causal inference, individualized treatment rules, matching, policy tree, regret bound
@@ 1 Introduction
The beneficial effects of treatments or interventions vary across individuals in most studies. For example, a training program that increases employment rate for some workers who would otherwise have remained unemployed, but others may remain unemployed despite intervention. Recently, data-driven personalized decision-making has received increased attention in randomized controlled trials and observational studies in various fields. In healthcare, clinicians utilize patients’ characteristics and medical history to tailor diabetes management to improve health outcomes (Bertsimas et al., 2017). In public policy, leveraging individual features lead to improved allocation of social services and thus enhance overall efficiency (Kube et al., 2019). In product recommendations, the online platforms effectively increase customer retention by recommending the products that match customers’ personal preferences (Bastani et al., 2022). A common underlying feature of these applications is the presence of treatment heterogeneity, which is a blessing that brings opportunities to improve the expected outcome in the target population. In view of substantially increasing volumes of personal data collected by advanced technologies nowadays, it is imperative to exploit heterogeneity of treatment effects efficiently and search for the optimal policy that allocates treatments based on individual characteristics.
@@
Statistical methods for learning the optimal policy have been extensively developed for more than a decade. They can be roughly divided into two categories: model-based methods and direct-search methods. Q-learning (Qian and Murphy, 2011) and A-learning (Shi et al., 2018) are model-based methods, which first fit the mean outcome given covariates and treatment or posit a model for the contrast function, and then determine the optimal policy based on the corresponding estimate. The performance of these methods critically relies on correct specification of the outcome model or the contrast function. Besides Q-learning, a long-standing common practice in policy learning has been estimating and maximizing the value function using weighting techniques. Outcome weighted learning (O-learning, Zhao et al., 2012) is a pioneering work of the direct-search methods that nonparametrically learns the optimal policy by utilizing inverse probability weighting (IPW) to estimate the value function in randomized clinical trials. In observational studies, the unknown propensity score in O-learning needs to be estimated using logistic regression or machine learning approaches (Lee et al., 2010). One recognized potential limitation of weighting-based approaches is the numerical instability caused by extreme values of the estimated propensity scores or model misspecification.
@@
To enhance efficiency and robustness of IPW-based learning methods, augmented inverse probability weighting (AIPW) has been introduced to policy learning (e.g., Zhang et al., 2012a; Zhang et al., 2012b; Zhao et al., 2019; Athey and Wager, 2021; Pan and Zhao, 2021). Zhang et al. (2012b) proposed to use AIPW to estimate the value function and showed that the estimation procedure is doubly robust in the sense that the estimator is consistent if either the parametric propensity score model or the parametric outcome mean model is correctly posited. However, the performance may be unsatisfactory if both models are misspecified (Kang and Schafer, 2007). Athey and Wager (2021) introduced AIPW-based policy learning using semiparametric theory, which leverages the cross-fitting technique (Chernozhukov et al., 2018) and flexible machine learning methods to nonparametrically estimate the nuisance functions: the propensity score and the conditional mean outcome. They proposed the policy tree algorithm to search for the optimal policy over a policy class. The learned policy is shown to achieve minimax optimal regret if the mean-squared errors of the nuisance functions satisfy certain order conditions of convergence rate. Despite the promising theoretical guarantee, the performance of AIPW-based learning deteriorates when extreme weights are present or when the sample size is small or moderate. Similar discussions can be found in the literature (e.g., Zhao et al., 2019; Wu et al., 2022). When the underlying conditional average treatment effect is not continuous at some points in the covariate space, Athey and Wager (2021) observed the interesting “phase transition” phenomenon, in which AIPW-based policy learning exhibited enormous variation in the estimated value function with the learned optimal policy when the sample size is moderate, whereas its performance is satisfactory for either small or large sample size.
Matching is widely used in many applied disciplines to infer causal effects (Imbens and Rubin, 2015; Rosenbaum, 2020). The goal of matching is to produce covariate balance such that the distributions of observed covariates in different treatment groups are approximately the same as they would be in a successful randomized experiment, thereby mimicking randomized experiments to the best extent with observational data (Stuart, 2010; Lin et al., 2023; Yang and Zhang, 2023). Therefore, matching has clearer intuitive appeal to practitioners compared to weighting techniques such as IPW and AIPW. Moreover, it is more robust to extreme values of the propensity score function in many circumstances (e.g., Wu et al., 2022; Yang and Zhang, 2023). Nevertheless, matching is rarely explored in policy learning except the work by Wu et al. (2020).
@@
In this paper, we develop a novel matching-based framework for policy learning in observational studies with favorable finite sample performance. The key idea behind the proposed matching-based policy learning (MB-learning) originates from an alternative form of the value function based on the pair of potential outcomes (referred to as advantage function in Section 2), which can be interpreted as the improvement on the expected outcome after implementing a given policy compared to a random policy with equal probabilities. We nonparametrically impute the missing potential outcomes using the response data within the matched set to develop a matching estimator of the advantage function, where we utilize nearest-neighbor (NN) matching on covariates with replacement (Abadie and Imbens, 2006). One recognized limitation of matching estimator is its susceptibility to the curse of dimensionality, and it is not $\sqrt{n}$-consistent due to a conditional bias term incurred by sampling with replacement during matching. To mitigate this issue, we employ the bias correction technique in Abadie and Imbens (2011) and develop an improved matching estimator of the advantage function with $\sqrt{n}$-consistency. Leveraging the formulation of the matching estimator, we recast the optimization problem of policy learning into a weighted classification problem and derive the optimal policy using the policy tree algorithm (Athey and Wager, 2021).
This paper makes the following major contributions to the existing literature. First, different from the (A)IPW-based approaches, MB-learning is a robust method for learning the optimal policy in observational studies. By utilizing a bias-corrected matching-based estimator, MB-learning mitigates the potential variability arising from propensity score model misspecification or extreme weights. Second, we establish theoretical guarantee for the optimal policy learned by the matching-based estimator, where a non-asymptotic high probability bound for the value function difference between the best-in-class policy and the learned policy is derived. Moreover, we demonstrate the superiority of MB-learning both theoretically and empirically. Athey and Wager (2021) established that AIPW-based policy learning with cross-fitting achieves rate-optimality in terms of minimax regret with the order of $\sqrt{\text{VC}(\Pi)/n}$, where $\text{VC}(\Pi)$ describes the complexity of a given policy class $\Pi$ which will be defined formally in the following sections. In comparison, we prove that the regret bound of MB-learning is almost of order $\sqrt{\text{VC}(\Pi)/n}$ for any fixed size of the matched set. Although MB-learning is not exactly rate-optimal, it exhibits more robust finite sample performance in simulation studies and real data analysis especially when the sample size is not large or when the policy allocations are extremely imbalanced. Moreover, the undesirable “phase transition” phenomenon encountered when implementing AIPW-based policy learning with cross-fitting does not occur when the proposed MB-learning is in use.
To the best of our knowledge, matching was rarely utilized for policy learning except matched learning (Wu et al., 2020, M-learning), where a general matched pair-based objective function is formulated and shown to be consistent of the value function for some special case and the optimal policy is learned by weighted support vector machine. We highlight several distinctive features of MB-learning compared to M-learning. First, rather than estimating the value function or its equivalent counterpart directly, the first step of M-learning is to construct a complicated matched pair-based objective function and then prove its consistency in theory. In contrast, MB-learning is more intuitive because it directly targets the advantage function and uses various matching techniques to obtain reasonable estimation. Second, we acknowledge and tackle the conditional bias due to matching, correlation due to sampling with replacement, and we fix the number of matches in the establishment of the asymptotic property and the non-asymptotic regret bound for MB-learning. In M-learning, there is a lack of discussions on correlation due to sampling with replacement, and the number of matches is forced to go to infinity with the sample size in theory. Third, the policy derived from MB-learning possesses appealing interpretability for practitioners because policy tree (Athey and Wager, 2021) is used in the optimization procedure, whereas M-learning with Gaussian kernels may be less interpretable. Finally, the bias introduced by NN matching may substantially deteriorate the performance of M-learning when moderate or high-dimensional covariates are present, whereas potential improvement using bias-corrected techniques remains unexplored in M-learning. In comparison, we mitigate this issue using the bias-corrected matching estimator (Abadie and Imbens, 2011) with Lasso in MB-learning.
@@
The rest of this paper is organized as follows. In Section 2, we develop the methodology of MB-learning. We first formulate the policy learning framework, introduce the advantage function, and describe AIPW-based policy learning developed by Athey and Wager (2021). Then, we propose a matching estimator of the advantage function. The estimator is further improved by a bias correction technique. In Section 3, we formulate MB-learning from a weighted classification perspective, and derive the optimal policy via policy tree using the bias-corrected matching estimator of the advantage function. We establish rigorous asymptotic and non-asymptotic theoretical guarantees for MB-learning. The empirical performance of MB-learning is demonstrated through extensive simulation studies in Section 4. An application to National Supported Work (NSW) Program is presented in Section 5. We conclude this article with discussions in Section 6. The proofs of the theories and additional numerical results are relegated to the Supplementary Material.
@@ 2.1 Notations and Preliminaries
Let $X\in\mathcal{X}$ be a vector of $p$-dimensional pre-treatment covariates, where $\mathcal{X}\subset\mathbb{R}^{p}$ is a covariate space, $W\in\{0,1\}$ is a binary policy or treatment indicator, and $Y$ is an outcome of interest. Without loss of generality, we assume that a larger $Y$ is more desirable. Under the potential outcome framework (Imbens and Rubin, 2015), we define $Y(w)$ to be the potential outcome if the unit had received the treatment $W=w$. We adopt the standard assumptions in causal inference: (i) consistency: $Y=Y(W)$; (ii) unconfoundedness: $\{Y(1),Y(0)\}\perp W\mid X$; and (iii) overlap: for any $x\in\mathcal{X}$, the propensity score $e(x)=\Pr(W=1\mid X=x)$ satisfies $c<e(x)<1-c$ for some positive constant $c$. Additionally, let $\mu(x,w)=E\left[Y|X=x,W=w\right]$, $\sigma^{2}(x,w)=\text{Var}\left[Y|X=x,W=w\right]$, and $\epsilon_{i}=Y_{i}-\mu(X_{i},W_{i})$. Under the assumptions of consistency and unconfoundedness, we have $\mu(x,w)=E\left[Y(w)|X=x\right]$ and $\sigma^{2}(x,w)=\text{Var}\left[Y(w)\mid X=x\right]$. Let $\tau=E[Y(1)-Y(0)]$ and $\tau(X)=E\left[Y(1)-Y(0)\mid X\right]$ represent the average treatment effect (ATE) and conditional average treatment effect (CATE), respectively. Furthermore, we assume that the observed data $\left\{(X_{i},W_{i},Y_{i})\right\}_{i=1}^{n}$ are $n$ independent and identically distributed (i.i.d.) copies of $(X,W,Y)$.
@@
We consider the problem of learning the optimal policy in observational studies, where we define that any policy $\pi$ is a deterministic decision function mapping from the covariate space $\mathcal{X}$ to the treatment set $\{0,1\}$. In practice, it is unrealistic to assign a complicated policy to the target population without restriction. One should attach problem-specific constraints to the policy, such as political restriction (Kitagawa and Tetenov, 2018), fairness (Athey and Wager, 2021) or interpretability (Zhang et al., 2012a). Therefore, we seek to learn the optimal policy within a restricted policy class $\Pi$.
The value function $V(\pi)$ for the policy $\pi$ (e.g., Zhao et al., 2012) is the expected outcome after implementing the treatment allocation by $\pi$, defined as $V(\pi)=E\left[Y(\pi(X))\right]$. The optimal policy $\pi^{*}$ within the class $\Pi$ maximizes the value function $V(\pi)$, i.e., $\pi^{*}=\arg\max_{\pi\in\Pi}V(\pi)$. The advantage function
$A(\pi)=E\left[\left(2\pi(X)-1\right)\tau(X)\right]$
@@
serves as an alternative form of the value function. Because $V(\pi)=E[Y(0)+(Y(1)-Y(0))\pi(X)]$, it follows that
$A(\pi)=2\left\{V(\pi)-\frac{E[Y(1)]+E[Y(0)]}{2}\right\}.$
@@
Hence, the advantage function $A(\pi)$ can be interpreted as twice the improvement achieved by implementing policy $\pi$ relative to a randomly assigned policy with equal probability (Mo et al., 2021). That is, it is a measure of evaluating the gain of implementing a given policy compared to the equiprobable random policy. The policy that maximizes $A(\pi)$ is equivalent to that maximizes $V(\pi)$. Throughout this paper, we focus on policy learning based on $A(\pi)$.
@@
By leveraging the idea of semiparametric efficient estimation (Robins et al., 1994), AIPW-based estimators of $V(\pi)$ or $A(\pi)$ were developed in the literature, and the optimal policy was determined based on these estimators (e.g., Zhang et al., 2012a; Athey and Wager, 2021). The AIPW-based estimator of the advantage function proposed by Athey and Wager (2021) is
$\hat{A}_{aipw}(\pi)=\frac{1}{n}\sum_{i=1}^{n}\left\{2\pi(X_{i})-1\right\}\Gamma_{i},$
where
$\Gamma_{i}=\mu(X_{i},1)-\mu(X_{i},0)+\frac{W_{i}-e(X_{i})}{e(X_{i})\left\{1-e(X_{i})\right\}}\left\{Y_{i}-\mu(X_{i},W_{i})\right\}.$
This AIPW-based estimator of $A(\pi)$ is efficient and doubly robust against model misspecification in observational studies. Athey and Wager (2021) employed black-box machine learning methods and the cross-fitting technique (Chernozhukov et al., 2018) to nonparametrically estimate the nuisance functions: the propensity score $e(X_{i})$ and the conditional mean outcome $\mu(X_{i},W_{i})$. The optimal policy is learned by maximizing the $\hat{A}_{aipw}(\pi)$ with the plugged-in estimated nuisance functions using the policy tree algorithm.
Nevertheless, the procedure of estimating the nuisance functions may introduce instability particularly when the estimated propensity score is extreme. In addition, machine learning methods may perform unsatisfactorily with a relatively small or moderate sample size. Weighting does not directly ensure covariate balance across treatment groups and is less intuitively appealing to practitioners as matching. Therefore, it is imperative to develop a sensible approach for learning the optimal policy with ensured covariate balance property and desirable efficiency and robustness performance in finite samples. Learning policies using matching techniques is a natural choice.